50 billion million deficit of electrons will have a charge of
1
Expert's answer
2021-09-14T09:37:14-0400

First we confirm how many electrons were removed:


"N=\text{50 billion million electrons}
\\N=(50)(10^9)(10^6)\text{ electrons}
\\N=5\times10^{16}\text{ electrons}"


Since we are removing electrons (because this is a deficit), the total charge will be positive:


"Q=5\times10^{16}\,\cancel {\text{electrons}}\times\dfrac{1.609\times10^{-19}\,C}{1\,\cancel {electron}}
\\ \text{ }
\\Q=8.045\times10^{-3}\,C=8.045\,mC"


In conclusion, 50 billion million deficit of electrons will have a charge of 8.045 X 10-3 C or 8.045 mC.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.